A rock is thrown straight vertically upward with a
Answer #1

9)

a)

initial velocity: V = 15 m/s

acceleration: g = 9.81 m/s^2

maximum height from throwing point = H = (V^2) / 2g = (15)^2 / (2 x 9.81) = 11.48m

b)

h = 11.48 m / 2 = 5.74 m

vf = vi + ad

vf = 15 + (9.81 x 5.74)

vf = 71.3

10)

the potential energy of a spring is 1/2 kx^2 where x is the distance stretched or compressed from equilibrium

3.6J = 1/2 k(0.12m)^2

k = (2 x 3.6J) / (0.12m)^2

k = 500 N/m

b) not sure..